Main Logbooks page LST commissioning Shift checklists Safe state checklists Maintenance activities LST2-4 Commissioning Old logbooks
GENERAL CAMERA DRIVE ENERGY AMC
  LST1 Drive ELOG, Page 2 of 3  ELOG logo
New entries since:Thu Jan 1 00:00:00 1970
ID Date Author Type Category Subject Status Text Attachments
  259   Tue Jul 2 21:25:07 2019 Dirk HoffmannApplicationsSoftwareCurrent & Power Monitoring stuck ("Socomec display") The Current&Power Monitoring was not
updated and hidden by two emtpy modal dialogue
windows.
  
  258   Fri Jun 28 23:06:36 2019 Dirk HoffmannTelescope OperationsGeneralTelescope stuck Ino explained: MOS was totally desynchronised
with PLC. (It was assuming that the telescope
was parked.) He parked the telescope manually.
  
  257   Fri Jun 28 22:49:39 2019 Dirk HoffmannTelescope OperationsGeneralTelescope stuck After tracking Mk421, we abort current
action (acknowledged) and try to go to a
new source: RA=308.1, Dec=40.96. Drive Control
  
  256   Thu Jun 27 12:19:02 2019 InoProblem FixedGeneralPLC Status During this weeks we solve this problems:

_ Park In: Sometimes the final
locking position was too limit and the "In
  
  Draft   Thu Jun 27 12:04:43 2019 InoProblem FixedGeneral  During this 2 weeks we solve this problems:

_ Park In: Sometimes the final
locking position was too limit and the "In
  
  254   Wed Jun 26 20:06:22 2019 InoPLC + Control Software debugging Tracking Trajectory We sent a tracking with with RA211,08 and
DEC29,52

1st time the goto position of azimuth
  
  253   Thu May 30 13:18:04 2019 AlessandroTelescope Operations semi expert gui (marginal) issues - During shutter tests, we were requested
to move the telescope many times between
two positions at different zenith. Sometimes,
  
  252   Fri Apr 26 23:17:07 2019 Alessandro for the drive teamTelescope OperationsIssuesdrive tracking issue 
We
encountered a couple of issue with tracking

- When we tried to track on polaris
  
  251   Thu Apr 18 21:21:25 2019 Drive Team- Ino&Thierry&AlessandroTelescope OperationsHardwareElevation limit switch issue 

The 2 swicthes are too far
 ElevationEndSwicth.png 
  250   Thu Apr 18 21:07:20 2019 Drive Team- Ino&Thierry&AlessandroTelescope Operations TESTS Parkout Parkin for fixing the bolt issue 
Different ParkOut and Parkin done
with different values for testing the torque
limitation. 

By decreasing
  
  249   Thu Apr 18 09:10:29 2019 InoPLC Software and ConfigurationHardwareOffset of the Zero Azimuth After pointing Polaris yesterday, it seems
that the zero position of the telescope was
incorrect but It's never had been calibrated
  
  248   Wed Apr 17 21:26:13 2019 Drive Team- Ino&Thierry&Alessandro, SatoshiTelescope OperationsGeneralTracking for mirrors alignment 
Tracking at position :

RA
= 101.288541 RA Offset = -0.6
DEC
= -16.71343 DEC Offset = 0.7


  
  247   Tue Apr 16 14:34:22 2019 Drive Team- InoTelescope OperationsHardwareCabling the azimuth end switches 
Cabling the azimuth end switches.
  
  246   Tue Apr 16 10:10:55 2019 Drive Team- Ino&ThierryPLC Software and Configuration Elevation setup 
1 degree has been added to the
Elevation accordingly to the last measurement
given by Olger
  
  245   Mon Apr 1 07:43:23 2019 Drive Team- Ino&Thierry&ArmandProblem FixedHardwareLST Parking control 
On Wenesday morning (27/03/19)
:

With Holger and Robert (MPI)
, many telescope "park in/park out"
  
  244   Mon Apr 1 07:36:04 2019 Drive Team- Ino&Thierry&ArmandTelescope OperationsGeneralTracking debug and setup 
Regulation of the motors (PID)
for slow and fast tracking and motor oscillations.
Implementation
and Debugging of the Semi-Expert control
  
  243   Thu Mar 28 10:34:41 2019 Drive Team- Ino&Thierry&Armand + Juli (IFAE)Telescope Operations High speed drive tests new session of high speed drive tests after
issues with Power Supply.


Tests of emergency stop
  
  242   Fri Mar 8 09:25:22 2019 Drive Team- Ino&Thierry&ArmandPLC + Control Software debuggingSoftwareTracking debug and setup 
Tracking procedure has been debugged
from the GUI-MOS (OPCUA Server with tracking
plugin)-PLC
Still some improvments to
  
  241   Fri Mar 8 09:20:29 2019 Drive Team- Ino&Thierry&ArmandTelescope OperationsGeneralHigh speed drive tests 
On Thrusday :

new
movements at 45% from -90 to 270 have been
done.
No possible movement at
  
  240   Wed Mar 6 16:30:56 2019 Drive Team- Ino&Thierry&ArmandPLC + Control Software debuggingSoftwareTracking mode debug 
Debugging of the MOS plugin corresponding
to the tracking 
  
ELOG V3.1.5-fc6679b